Jocelynnn:

If 20 articles cost ? 90, then the cost of 9 articles is ?

Tonycoolkid21:

One way is: Find the unit rate first (or the cost of 1 article) . Then multiply the unit rate by 9 to find the cost of 9 articles. $90 = x 20 articles 1 article $90 ÷20 = $4.50 20 articles ÷ 20 1 article $4.50 • 9 = $40.50 for 9 articles

Tonycoolkid21:

Cost of 20 articles = 90 Rs Cost of 1 article = 90 / 20 = 4.5 Rs Cost of 9 articles = 9 X 4.5 = 40.5 Rs

Tonycoolkid21:

did i do good?

Jocelynnn:

@tonycoolkid21 wrote:
did i do good?
yes u did lol

Tonycoolkid21:

medal me i helped you